a. Calculate the pH of 0.100 L of a buffer solution that is 0.29M in HF (Ka = 3.5 x 10-4 ) and 0.55M in NaF.

- What is the pH after adding 0.004mol of HNO3 to the buffer described in Part A

-What is the pH after adding 0.002mol of KOH to the buffer described in Part A?
Answer #1

We know, Moles = Molarity (M) x Volume (V in L) So Moles HF = 0.2900 * 0.100 mol = 0.0290 mol Moles F- = 0.5500 * 0.100 = 0.0